A delivery company accepts only rectangular boxes the sum of whose length and the perimeter of a cross-section does not exceed 108 inches. Find the dimensions of an acceptable box of largest volume. 


1
Expert's answer
2022-01-10T15:27:55-0500

L+4S=108

L=108-4S

"v=s^2L\\ v=s^2(108-4s)\\
v=108s^2-4s^3\\
v'=216s-12s^2\\ 0=216s-12s^2\\
s=0,s=18\\
then, L=108-4(18)\\
L=36.\\
v=s^2L, v= 18^2(36),\\
v=11664in^3"